The Volume of a cube is given by V=s^3, where s is the length of a side. Find the length of a side of acube if the volume is 1000 cm^3
Answer by Nate(3500)

V = s^3
1000 = s^3
(1000)^(1/3) = s = 10
Each side of the cube is 10 cm.